OpenStudy (anonymous):

combine {sqrt -8} + {sqrt -32}

OpenStudy (anonymous):

The answer in my book is \[6i - 2i \sqrt{10}\] I dont know how they got that answer.

OpenStudy (anonymous):

You can't take the square root of a negative number

OpenStudy (anonymous):

@ksaimouli is incorrect

OpenStudy (anonymous):

thats why you multiply it by i which is equal to -1

OpenStudy (anonymous):

@ksaimouli That's totally wrong. sqrt(a)+sqrt(b) is NOT sqrt(a+b). @ilikephysics2 Yes, You can. That's what imaginary numbers are there for. @karinewoods17 Notice that \[ \Large {\sqrt {8} = 2\sqrt{2}} \]and \[ \Large {\sqrt {32} = 4\sqrt{2}}, \]so \[ \Large {\sqrt {-8} = 2i\sqrt{2}} \]and \[ \Large {\sqrt {-32} = 4i\sqrt{2}} \]Do you see how that works? Now, just add to get 6isqrt(2).

OpenStudy (anonymous):

wouldnt it be \[6i \sqrt{2}\] ??

OpenStudy (anonymous):

\[\sqrt{-8}+\sqrt{-32}=\sqrt{-8}+\sqrt{-8\times4}= \sqrt{-8} + 2\sqrt{-8}\]

OpenStudy (anonymous):

\[3\sqrt{-8}=3i \sqrt{8}=3i \sqrt{2\times4}=6i \sqrt{2}\]

OpenStudy (anonymous):

Yes you are correct

OpenStudy (anonymous):

Yep. That's exactly what I said, except I explained everything. =P

OpenStudy (anonymous):

but... the answer in my book says the answer is \[6i - 2i \sqrt{10}\] How did they get that?

OpenStudy (anonymous):

Im sorry:/ Im making coffee and waiting for you:)

OpenStudy (anonymous):

im not! sorry!

OpenStudy (anonymous):

I was making breakfast for my boyfriend.

OpenStudy (anonymous):

umm. \[\sqrt{-32}\] breaks into \[\sqrt{-32} = \sqrt{4} * \sqrt{-8}\] which simplifies to \[4i \sqrt{2}\] ?

OpenStudy (anonymous):

And yes I try:)

OpenStudy (anonymous):

now what. lol

OpenStudy (anonymous):

@nincompoop please do not ignore me;)